OpenStudy (anonymous):

How to I factor n^5+216n^2 ?

OpenStudy (b77w):

Jiggly nice!

OpenStudy (b77w):

n^2(n+6)(n^2-6n+36)

OpenStudy (anonymous):

How do I get that?

OpenStudy (shamim):

n^2(n^3+216) n^2(n^3+6^3) =n^2(n+6)(n^2-6n+6^2)

OpenStudy (cwrw238):

n^2 in the greatest common factor of n^5 and 216n^2 so that can be taken out to give n^2(n^3 + 216) the brackets contain 2 perfect cubes and are factored using the identity a^3 + b^3 = (a + b)(a^2 - ab + b^2)

OpenStudy (anonymous):

Thank you to everyone! I understand now.

OpenStudy (shamim):

welcome